A)To find the level of measurement : nominal,ratio,interval and ordinal


1)Jason measures time on days, with 0 corresponding to his birth date. The day before his birth is -1, the day after his birth date is +1, and so on. Jason has converted dates of major historical events to his numbering systems. What is the level of measurement of these numbers?


2)Which is relatively better: A score of 85 on a chemistry test or a score of 45 on an physics test? Scores on the chemistry test have a mean of 90 and a standard deviation of 10. Scores on physics test have a mean of 55 and a standard deviation of 5


1
Expert's answer
2021-01-14T20:03:54-0500

1) Ratio level.


2) For chemistry: "z=\frac{85-90}{10}=-0.5."

For physics: "z=\frac{45-55}{5}=-2."

Score on a chemistry is relatively better.